OpenStudy (undeaddragon26):

Find the radius of a cone with volume 45.19 cm3 and height 7.5 cm. Answer 1.2 cm 2.4 cm 2.6 cm 4.8 cm

OpenStudy (anonymous):

@AkashdeepDeb are you gonna reply yet

OpenStudy (akashdeepdeb):

Volume of a cone = \[\frac{1}{3} \pi r^2 h = 45.19~~~~~~[Given]\] \[r = 7.5 cm\] \[\frac{1}{3} \pi (7.5)^2 h = 45.19\] Now can you find 'h' ? HINT: Re-arrange the equation. Are you getting this? :)